XY2 dissociates as XY2 (g) gives XY(g)+Y(g) When the initial pressure of XY2 is 600mmHg ,the total equilibrium pressure is 800mmHg.calculate Kp for the reaction assuming that the volume of the system remains unchanged.

XY2  dissociates as
XY2 (g) gives XY(g)+Y(g)
When the initial pressure of XY2 is 600mmHg ,the total equilibrium pressure is 800mmHg.calculate Kp for the reaction assuming that the volume of the system remains unchanged.

Grade:12th pass

1 Answers

Arun
5 years ago
Initial pressure  a = 600 mm Hg 
                      XY2 (g) -------------------------->  XY(g)  +  Y(g)
 at equilibrium  a - x                                     x            x
at equilibrium total pressure  ( a - x + x + x)  =800 mm Hg
                                           600 + x = 800
                                               x = 200
 Kp  =  Pxy * Py / Pxy2
        =   200 * 200 / 400
        =  100 mm Hg 
        =   100 / 760 atm
        =   0.132 atm